OpenStudy (anonymous):

how do i solve this for x? x^2=1/2 and x^2=3/2

OpenStudy (anonymous):

square root both sides.

OpenStudy (anonymous):

right, but i don't understand how to do that. the example given is this: x^2=2/3 and x^2=3/5 and somehow they get this: x=\[\pm \sqrt{6}/3 and \pm \sqrt{15}/5\]

OpenStudy (anonymous):

i don't understand the steps between the two

OpenStudy (anonymous):

x=+sqrt(1/2) or x+-sqrt(1/2)

OpenStudy (anonymous):

the same for 3/2